The difference between the T distribution and the standard normal (z) distribution decreases as the degrees of freedom increase.

True

False

When estimating the number of responses to be collected in order to achieve a desired margin of error on a proportion, what value should be used for the population proportion if no estimate is available? Enter your answer as a decimal to 2 decimal places.

A population is believed to have a standard deviation of 23.1. An analyst wants to create a 95% confidence interval with a maximum margin of error of 2.6. What is the minimum sample size required to achieve this goal? Enter a whole number only.

A population is believed to have a standard deviation of 13. A sample of size 96 is taken and the sample mean is calculated as 768.9. What is the margin of error (E) for a 95% confidence interval for the mean? Enter your answer to 2 decimal places.
Answer #1

Ans:

1)True,as the degree of freedom increases,shape of the t distribution approches to the normal distribution.

2)When no prior estimate is available,we use p=0.50

3)minimum sample size required,n=(1.96*23.1/2.6)^2=303

4)Margin of error=1.96*(13/sqrt(96))=2.60